# Collaborating Vision, Depth, and Thermal Signals for Multi-Modal Tracking: Dataset and Algorithm

### v2

@@ 1 Introduction
Visual object tracking aims to automatically localise an object of interest within a video, based on the initially specified object position and scale (javed2023visual; liu2024spatial). It is a fundamental research topic in the field of artificial intelligence and computer vision. Through decades of research, visual object tracking has witnessed substantial progress, driven by continuous advancements in benchmark datasets (huang2019got; fan2019lasot; zhu2024unimod1k) and tracking algorithms (xu2021adaptive; xu2023learning; zhu2022robust). However, conventional tracking methods mainly rely on visible RGB images, often suffer from reduced effectiveness and robustness under conditions of adverse visibility (tang2025revisiting).
To cope with these challenges, recent research has explored the integration of an additional sensing modality to enhance tracking performance in complex and visually degraded scenarios. For instance, multi-modal tracking methods, such as RGB-D and RGB-T tracking, leverage complementary information from depth and thermal data, respectively, to improve robustness and accuracy in adverse conditions. In particular, RGB-D tracking (zhu2024self; zhu2023rgbd1k) combines an RGB image with a depth map to provide additional spatial and structural cues, enabling more accurate localisation in challenging scenarios such as occlusion, background clutter, and scale variation. Similarly, RGB-T tracking (tang2024generative; xiao2022attribute) integrates RGB and thermal infrared images, compensating for the limitations of visible light in low-illumination scenarios.
@@
In detail, to leverage the representation power of the pre-trained RGB tracking model and effectively integrate tri-modal information for object localisation, RDTTrack incorporates a prompt learning mechanism. Specifically, to fuse depth and thermal infrared cues, it utilises a feature projection to enforce orthogonality between depth and thermal infrared features. This orthogonal projection aims to reduce feature redundancy and enhance the reliability of the fused representation. Subsequently, fused depth and thermal features are integrated with RGB features as prompts for the pre-trained OSTrack (ye2022joint) model, effectively leveraging tri-modal complementary cues for robust tracking. For RDTTrack training, the pre-trained OSTrack model is kept frozen, and only the prompt learning module is fine-tuned using the 400 training sequences from the RGBDT500 dataset. Extensive experiments, including ablation studies and comparative evaluations, demonstrate the effectiveness and the competitive performance of the proposed baseline tracker.
@@ 2.1 Multi-Modal Tracking Datasets
Developing multi-modal tracking datasets is essential for both advancing and rigorously evaluating tracking algorithms. To date, most existing datasets have concentrated on RGB-D and RGB-T modalities. Among RGB-D tracking datasets, DepthTrack (yan2021depthtrack) stands out as a significant benchmark, comprising 200 video sequences. More recently, the RGBD1K dataset zhu2023rgbd1k has been introduced, featuring 1,050 sequences, with object bounding box annotations provided for around 720K RGB-D frames. These large-scale datasets have been instrumental in advancing the development of RGB-D tracking approaches. Additionally, other important benchmarks, such as CDTB lukezic2019cdtb, PTBsong2013tracking, and STC xiao2017robust, are also valuable for evaluating RGB-D tracking algorithms.
In RGB-T tracking, several influential datasets have been introduced, including GTOT li2016learning, RGBT234 li2019rgb, and LasHeR li2021lasher. GTOT serves as the first benchmark dataset for RGB-T tracking, comprising 50 RGB-T video sequences. Building upon the GTOT benchmark, RGBT234 extends the scale to 234 RGB-T sequences, providing broader coverage of diverse tracking scenarios. To date, LasHeR represents the most extensive RGB-T tracking dataset, containing 1,244 RGB-T sequences and over 730K paired frames. The thermal modality, which is particularly effective under low-light and adverse illumination conditions, enhances tracking robustness in real-world environments. The release of these datasets has significantly contributed to the advancement of RGB-T tracking.
@@ 2.2 Multi-Modal Tracking Methodologies
RGB-D tracking leverages complementary information from both RGB and depth modalities, enabling more comprehensive scene understanding for tracking. A variety of trackers have been proposed to realise effective fusion and interaction between RGB and depth modalities. For example, DAL (qian2021dal) integrates depth information into RGB features based on the correlation filter-based tracking framework. Recent trackers, such as DeT (yan2021depthtrack), adapt RGB-only tracking architectures (danelljan2019atom; bhat2019learning) by fusing feature maps through pixel-wise operations. SPT (zhu2023rgbd1k) utilises separated transformer encoders for RGB and depth, followed by a dedicated fusion module. ARKitTrack (zhao2023arkittrack) further advances fusion by encoding depth into bird’s eye view representations and performing cross-view fusion. Collectively, these approaches represent ongoing progress in RGB-D tracking, aiming to enhance multi-modal feature extraction, fusion, and adaptation for improved tracking robustness.
@@
RGB-T tracking leverages the complementary strengths of RGB and thermal modalities to enhance tracking performance under challenging conditions such as low-light, nighttime environments. Early efforts primarily focus on convolutional neural network-based fusion strategies (tang2025omni; he2016deep). For example, mfDiMP (zhang2019multi) introduces an end-to-end framework for RGB-T fusion, while MANet (li2019multi) employes a three-way adapter network to extract modality-specific and shared features between two modalities. APFNet (xiao2022attribute) introduces an adaptive feature fusion mechanism that fine-tunes attributes between different modalities, thus enhancing the robustness and accuracy of the algorithm. Furthermore, with the advent of Vision Transformers (dosovitskiy2020image), RGB-T tracking technology has seen further innovations. For example, by introducing visual prompt learning zhu2023visual; wu2024single; wang2024temporal; hong2024onetracker into the Transformer-based tracking frameworks, auxiliary thermal information is fused into pre-trained RGB-only models, significantly improving tracking performance. Additionally, BAT cao2024bi and TBSI hui2023bridging propose feature bridging mechanisms to strengthen cross-modal interactions.
@@
It is also noteworthy that research on pixel-level RGB-T image fusion provides valuable guidance for cross-modal information integration in tracking (liu2024promptfusion). The comprehensive review by (liu2024infrared), which systematically explores the field from data compatibility to task adaptation, offers a solid theoretical foundation for understanding how to effectively fuse pixel-level heterogeneous modalities to serve high-level vision tasks (jiang2024multispectral; liu2022attention). For instance, CoCoNet (liu2024coconet) introduces a multi-level feature ensemble for multi-modal image fusion, significantly advancing the performance in both image fusion and downstream object detection tasks. Through pixel-level fusion, complementary information from RGB and TIR modalities can be effectively aligned and enhanced (song2025refinefuse), thereby providing a more robust and information-rich joint representation for subsequent trackers.
@@
UAV123 benchmark2016benchmark 123 113K - ECCV 2016
GOT-10k huang2019got 10K 1.5M 563 IEEE TPAMI 2019
LaSOT fan2019lasot 1400 3.5M 70 CVPR 2019
GTOT li2016learning 50 15.8K 9 IEEE TIP 2016
RGBT210 li2017weighted 210 104.7K 22 ACM MM 2017
RGBT234 li2019rgb 234 116.7K 22 PR 2019
LasHeR li2021lasher 1224 734.8K 32 IEEE TIP 2021
VTUAV zhang2022visible 500 1.7M 13 CVPR 2022
PTB song2013tracking 100 21K 26 CVPR 2013
CDTB lukezic2019cdtb 80 101.9K 21 ICCV 2019
DepthTrack yan2021depthtrack 200 294.5K 90 ICCV 2021
ARKitTrack zhao2023arkittrack 455 229.7K 144 CVPR 2023
@@ 3.1 Dataset Details
The RGBDT500 dataset includes a diverse range of object categories, covering over 66 classes, including household items, animals, and vehicles, as shown in Fig. 2. In constructing the categories, we strategically consider the unique strengths and limitations of each modality. For example, sequences featuring objects in low-light conditions or containing high-temperature objects are deliberately captured to emphasise the advantages of the thermal modality. Conversely, we capture several video sequences featuring planar objects with minimal geometric depth variation to intentionally limit the contribution of the depth modality, thereby encouraging greater reliance on RGB and thermal infrared cues. To further enhance the dataset’s complexity and realism, we capture some sequences exhibiting modal perturbations, such as inconsistent viewpoints among RGB, depth, and thermal infrared streams. Collectively, these design choices make RGBDT500 a challenging and comprehensive benchmark for evaluating the robustness of tri-modal tracking algorithms. In addition, to preserve privacy, we have applied EgoBlur raina2023egoblur model to blur all recognizable faces and license plates present in the images.
@@ 3.2 Data Annotation
The RGBDT500 is annotated with accurate object bounding boxes for frames containing the object of interest, providing reliable ground truth for developing and assessing tracking algorithms. For the training set, to minimise annotation costs and while meeting requirements for training, we employ K-means clustering (hartigan1979algorithm) to select the most representative frames from each training sequence, which are subsequently annotated with object bounding boxes. Besides, the test set includes 100 sequences, encompassing 43.7K RGB-D-T image triplets with dense bounding box annotations, thereby facilitating rigorous and detailed performance assessment. For the bounding box annotation, we adopt the top-left corner coordinates $(x,y)$, along with the width $w$ and height $h$ of the target’s bounding box, to represent the ground truth in the format $[x,y,w,h]$.
@@ 3.3 Evaluation Metrics
For evaluation, we follow the One-Pass Evaluation (OPE) protocol (wu2015object) to assess tracking performance on RGBDT500. Specifically, we compute the Distance Precision (DP) and the Area Under the Curve (AUC) of the success plot to quantitatively measure the effectiveness of multi-modal tracking methods. Specifically, the precision is measured by computing the distance between the predicted and the ground-truth bounding boxes. By varying a predefined distance threshold to determine successful tracking, a precision plot can be generated to illustrate tracker performance. Based on the precision plot, DP is defined as the percentage of frames in which the prediction error falls within a threshold of 20 pixels. In addition, tracking success can also be evaluated by determining whether the overlap ratio between the predicted and ground-truth bounding boxes exceeds a predefined threshold. The success plot illustrates the proportion of successful frames across a range of overlap thresholds from 0 to 1. Then, the AUC of the success plot is computed to quantitatively assess trackers.
@@ 4 The Tri-Modal Tracking Baseline
To support tri-modal object tracking on RGBDT500 and promote progress toward more general multi-modal tracking, we design a specialised tracking algorithm, RDTTrack, capable of effectively integrating RGB, depth, and thermal infrared information. The RDTTrack is developed by extending a pre-trained RGB-only tracker, OSTrack ye2022joint, with a specially designed prompt learning module that enables integration of tri-modal cues for tracking.
@@ 4.1 Overall Architecture
For the training of RDTTrack, we freeze all parameters of the RGB streams, and fine-tune only the Depth-TIR fusion module and all prompt learning blocks. The overall loss function of RDTTrack is a combination of focal loss lin2017focal for classification, $L_{1}$ loss and the GIoU loss rezatofighi2019generalized for localisation, calculated as: $\mathcal{L}=L_{\text{CLS}}+\lambda_{\text{GIoU}}\cdot L_{\text{GIoU}}+\lambda_{L_{1}}\cdot L_{1}$, where $\lambda_{\text{GIoU}}$ and $\lambda_{L_{1}}$ are two constants.
@@ 4.3 Multi-Modal Prompt Learning Blocks
Fig. 3(c) presents the architecture of the multi-modal prompt learning module. It is designed to extract complementary information between the RGB modality and the fused auxiliary D-TIR modalities to produce more informative visual prompts. The prompt learning block follows the design proposed in (zhu2023visual). The $l$-th prompt learning block $\mathrm{Propmt}^{l}()$ takes as input the tokens $\bm{H}^{l-1}$ from (l1)(l-1)-th Transformer encoder layer along with the multi-modal prompts $\bm{P}^{l-1}$ produced by the preceding prompt block, as:
@@
$\odot$ denotes element-wise multiplication. For the first prompt learning block, its input is the RGB tokens $\bm{T}_{\mathrm{RGB}}$ and the D-TIR fused representation $\bm{T}_{\mathrm{D-TIR}}$. For more detailed information of the prompt learning block, readers are referred to the reference (zhu2023visual).
@@ 5.1 Experimental Settings
ATOM danelljan2019atom RGB 0.649 0.695 2019
DiMP bhat2019learning RGB 0.662 0.710 2019
PrDiMP danelljan2020probabilistic RGB 0.676 0.698 2020
ToMP mayer2022transforming RGB 0.710 0.747 2022
OSTrack ye2022joint RGB 0.694 0.736 2022
SeqTrack chen2023seqtrack RGB 0.719 0.766 2023
MixFormer cui2024mixformer RGB 0.732 0.781 2024
DeT_ATOM yan2021depthtrack RGB+D 0.639 0.665 2021
DeT_DiMP yan2021depthtrack RGB+D 0.667 0.700 2021
SPT zhu2023rgbd1k RGB+D 0.706 0.757 2023
ViPT_RGBD zhu2023visual RGB+D 0.720 0.759 2023
SDSTrack_RGBD hou2024sdstrack RGB+D 0.718 0.763 2024
UnTrack_RGBD wu2024single RGB+D 0.733 0.776 2024
TBSI hui2023bridging RGB+T 0.690 0.749 2023
ViPT_RGBT zhu2023visual RGB+T 0.693 0.752 2023
SDSTrack_RGBT hou2024sdstrack RGB+T 0.666 0.708 2024
BAT cao2024bi RGB+T 0.713 0.782 2024
UnTrack_RGBT wu2024single RGB+T 0.732 0.790 2024
DCEvo+OSTrack liu2025dcevo RGB+T 0.706 0.741 2025

### v3

@@ 1 Introduction
Visual object tracking aims to automatically localise an object of interest within a video, based on the initially specified object position and scale (Javed et al., 2023; Liu et al., 2024a; Kristan et al., 2023). It is a fundamental research topic in the field of artificial intelligence and computer vision. Through decades of research, visual object tracking has witnessed substantial progress, driven by continuous advancements in benchmark datasets (Huang et al., 2019; Fan et al., 2019; Zhu et al., 2024b) and tracking algorithms (Xu et al., 2021, 2023; Zhu et al., 2022; Xu et al., 2020). However, conventional tracking methods mainly rely on visible RGB images, often suffer from reduced effectiveness and robustness under conditions of adverse visibility (Tang et al., 2025a).
To cope with these challenges, recent research has explored the integration of an additional sensing modality to enhance tracking performance in complex and visually degraded scenarios. For instance, multi-modal tracking methods, such as RGB-D and RGB-T tracking, leverage complementary information from depth and thermal data, respectively, to improve robustness and accuracy in adverse conditions. In particular, RGB-D tracking (Zhu et al., 2024a, 2023b) combines an RGB image with a depth map to provide additional spatial and structural cues, enabling more accurate localisation in challenging scenarios such as occlusion, background clutter, and scale variation. Similarly, RGB-T tracking (Tang et al., 2024; Xiao et al., 2022) integrates RGB and thermal infrared images, compensating for the limitations of visible light in low-illumination scenarios.
@@
In detail, to leverage the representation power of the pre-trained RGB tracking model and effectively integrate tri-modal information for object localisation, RDTTrack incorporates a prompt learning mechanism. Specifically, to fuse depth and thermal infrared cues, it utilises a feature projection to enforce orthogonality between depth and thermal infrared features. This orthogonal projection aims to reduce feature redundancy and enhance the reliability of the fused representation. Subsequently, fused depth and thermal features are integrated with RGB features as prompts for the pre-trained OSTrack (Ye et al., 2022) model, effectively leveraging tri-modal complementary cues for robust tracking. For RDTTrack training, the pre-trained OSTrack model is kept frozen, and only the prompt learning module is fine-tuned using the 400 training sequences from the RGBDT500 dataset. Extensive experiments, including ablation studies and comparative evaluations, demonstrate the effectiveness and the competitive performance of the proposed baseline tracker.
@@ 2.1 Multi-Modal Tracking Datasets
Developing multi-modal tracking datasets is essential for both advancing and rigorously evaluating tracking algorithms. To date, most existing datasets have concentrated on RGB-D and RGB-T modalities. Among RGB-D tracking datasets, DepthTrack (Yan et al., 2021b) stands out as a significant benchmark, comprising 200 video sequences. More recently, the RGBD1K dataset Zhu et al. (2023b) has been introduced, featuring 1,050 sequences, with object bounding box annotations provided for around 720K RGB-D frames. These large-scale datasets have been instrumental in advancing the development of RGB-D tracking approaches. Additionally, other important benchmarks, such as CDTB Lukezic et al. (2019), PTBSong and Xiao (2013), and STC Xiao et al. (2017), are also valuable for evaluating RGB-D tracking algorithms.
In RGB-T tracking, several influential datasets have been introduced, including GTOT Li et al. (2016), RGBT234 Li et al. (2019b), and LasHeR Li et al. (2021). GTOT serves as the first benchmark dataset for RGB-T tracking, comprising 50 RGB-T video sequences. Building upon the GTOT benchmark, RGBT234 extends the scale to 234 RGB-T sequences, providing broader coverage of diverse tracking scenarios. To date, LasHeR represents the most extensive RGB-T tracking dataset, containing 1,244 RGB-T sequences and over 730K paired frames. The thermal modality, which is particularly effective under low-light and adverse illumination conditions, enhances tracking robustness in real-world environments. The release of these datasets has significantly contributed to the advancement of RGB-T tracking.
@@ 2.2 Multi-Modal Tracking Methodologies
RGB-D tracking leverages complementary information from both RGB and depth modalities, enabling more comprehensive scene understanding for tracking. A variety of trackers have been proposed to realise effective fusion and interaction between RGB and depth modalities. For example, DAL (Qian et al., 2021) integrates depth information into RGB features based on the correlation filter-based tracking framework. Recent trackers, such as DeT (Yan et al., 2021b), adapt RGB-only tracking architectures (Danelljan et al., 2019; Bhat et al., 2019) by fusing feature maps through pixel-wise operations. SPT (Zhu et al., 2023b) utilises separated transformer encoders for RGB and depth, followed by a dedicated fusion module. ARKitTrack (Zhao et al., 2023) further advances fusion by encoding depth into bird’s eye view representations and performing cross-view fusion. Collectively, these approaches represent ongoing progress in RGB-D tracking, aiming to enhance multi-modal feature extraction, fusion, and adaptation for improved tracking robustness.
@@
RGB-T tracking leverages the complementary strengths of RGB and thermal modalities to enhance tracking performance under challenging conditions such as low-light, nighttime environments. Early efforts primarily focus on convolutional neural network-based fusion strategies (Tang et al., 2025b; He et al., 2016). For example, mfDiMP (Zhang et al., 2019) introduces an end-to-end framework for RGB-T fusion, while MANet (Li et al., 2019a) employes a three-way adapter network to extract modality-specific and shared features between two modalities. APFNet (Xiao et al., 2022) introduces an adaptive feature fusion mechanism that fine-tunes attributes between different modalities, thus enhancing the robustness and accuracy of the algorithm. Furthermore, with the advent of Vision Transformers (Dosovitskiy, 2020), RGB-T tracking technology has seen further innovations. For example, by introducing visual prompt learning Zhu et al. (2023a); Wu et al. (2024); Wang et al. (2024); Hong et al. (2024) into the Transformer-based tracking frameworks, auxiliary thermal information is fused into pre-trained RGB-only models, significantly improving tracking performance. Additionally, BAT Cao et al. (2024) and TBSI Hui et al. (2023) propose feature bridging mechanisms to strengthen cross-modal interactions.
@@
It is also noteworthy that research on pixel-level RGB-T image fusion provides valuable guidance for cross-modal information integration in tracking (Liu et al., 2024b; Cheng et al., 2025). The comprehensive review by (Liu et al., 2024d), which systematically explores the field from data compatibility to task adaptation, offers a solid theoretical foundation for understanding how to effectively fuse pixel-level heterogeneous modalities to serve high-level vision tasks (Jiang et al., 2024; Liu et al., 2022). For instance, CoCoNet (Liu et al., 2024c) introduces a multi-level feature ensemble for multi-modal image fusion, significantly advancing the performance in both image fusion and downstream object detection tasks. Through pixel-level fusion, complementary information from RGB and TIR modalities can be effectively aligned and enhanced (Song et al., 2025), thereby providing a more robust and information-rich joint representation for subsequent trackers.
@@
UAV123 Benchmark (2016) 123 113K - ECCV 2016
GOT-10k Huang et al. (2019) 10K 1.5M 563 IEEE TPAMI 2019
LaSOT Fan et al. (2019) 1400 3.5M 70 CVPR 2019
GTOT Li et al. (2016) 50 15.8K 9 IEEE TIP 2016
RGBT210 Li et al. (2017) 210 104.7K 22 ACM MM 2017
RGBT234 Li et al. (2019b) 234 116.7K 22 PR 2019
LasHeR Li et al. (2021) 1224 734.8K 32 IEEE TIP 2021
VTUAV Zhang et al. (2022) 500 1.7M 13 CVPR 2022
PTB Song and Xiao (2013) 100 21K 26 CVPR 2013
CDTB Lukezic et al. (2019) 80 101.9K 21 ICCV 2019
DepthTrack Yan et al. (2021b) 200 294.5K 90 ICCV 2021
ARKitTrack Zhao et al. (2023) 455 229.7K 144 CVPR 2023
@@ 3.1 Dataset Details
The RGBDT500 dataset includes a diverse range of object categories, covering over 66 classes, including household items, animals, and vehicles, as shown in Fig. 2. In constructing the categories, we strategically consider the unique strengths and limitations of each modality. For example, sequences featuring objects in low-light conditions or containing high-temperature objects are deliberately captured to emphasise the advantages of the thermal modality. Conversely, we capture several video sequences featuring planar objects with minimal geometric depth variation to intentionally limit the contribution of the depth modality, thereby encouraging greater reliance on RGB and thermal infrared cues. To further enhance the dataset’s complexity and realism, we capture some sequences exhibiting modal perturbations, such as inconsistent viewpoints among RGB, depth, and thermal infrared streams. Collectively, these design choices make RGBDT500 a challenging and comprehensive benchmark for evaluating the robustness of tri-modal tracking algorithms. In addition, to preserve privacy, we have applied EgoBlur Raina et al. (2023) model to blur all recognizable faces and license plates present in the images.
@@ 3.2 Data Annotation
The RGBDT500 is annotated with accurate object bounding boxes for frames containing the object of interest, providing reliable ground truth for developing and assessing tracking algorithms. For the training set, to minimise annotation costs and while meeting requirements for training, we employ K-means clustering (Hartigan and Wong, 1979) to select the most representative frames from each training sequence, which are subsequently annotated with object bounding boxes. Besides, the test set includes 100 sequences, encompassing 43.7K RGB-D-T image triplets with dense bounding box annotations, thereby facilitating rigorous and detailed performance assessment. For the bounding box annotation, we adopt the top-left corner coordinates $(x,y)$, along with the width $w$ and height $h$ of the target’s bounding box, to represent the ground truth in the format $[x,y,w,h]$.
@@ 3.3 Evaluation Metrics
For evaluation, we follow the One-Pass Evaluation (OPE) protocol (Wu et al., 2015) to assess tracking performance on RGBDT500. Specifically, we compute the Distance Precision (DP) and the Area Under the Curve (AUC) of the success plot to quantitatively measure the effectiveness of multi-modal tracking methods. Specifically, the precision is measured by computing the distance between the predicted and the ground-truth bounding boxes. By varying a predefined distance threshold to determine successful tracking, a precision plot can be generated to illustrate tracker performance. Based on the precision plot, DP is defined as the percentage of frames in which the prediction error falls within a threshold of 20 pixels. In addition, tracking success can also be evaluated by determining whether the overlap ratio between the predicted and ground-truth bounding boxes exceeds a predefined threshold. The success plot illustrates the proportion of successful frames across a range of overlap thresholds from 0 to 1. Then, the AUC of the success plot is computed to quantitatively assess trackers.
@@ 4 The Tri-Modal Tracking Baseline
To support tri-modal object tracking on RGBDT500 and promote progress toward more general multi-modal tracking, we design a specialised tracking algorithm, RDTTrack, capable of effectively integrating RGB, depth, and thermal infrared information. The RDTTrack is developed by extending a pre-trained RGB-only tracker, OSTrack Ye et al. (2022), with a specially designed prompt learning module that enables integration of tri-modal cues for tracking.
@@ 4.1 Overall Architecture
For the training of RDTTrack, we freeze all parameters of the RGB streams, and fine-tune only the Depth-TIR fusion module and all prompt learning blocks. The overall loss function of RDTTrack is a combination of focal loss Lin (2017) for classification, $L_{1}$ loss and the GIoU loss Rezatofighi et al. (2019) for localisation, calculated as: $\mathcal{L}=L_{\text{CLS}}+\lambda_{\text{GIoU}}\cdot L_{\text{GIoU}}+\lambda_{L_{1}}\cdot L_{1}$, where $\lambda_{\text{GIoU}}$ and $\lambda_{L_{1}}$ are two constants.
@@ 4.3 Multi-Modal Prompt Learning Blocks
Fig. 3(c) presents the architecture of the multi-modal prompt learning module. It is designed to extract complementary information between the RGB modality and the fused auxiliary D-TIR modalities to produce more informative visual prompts. The prompt learning block follows the design proposed in (Zhu et al., 2023a). The $l$-th prompt learning block $\mathrm{Propmt}^{l}()$ takes as input the tokens $\bm{H}^{l-1}$ from (l1)(l-1)-th Transformer encoder layer along with the multi-modal prompts $\bm{P}^{l-1}$ produced by the preceding prompt block, as:
@@
$\odot$ denotes element-wise multiplication. For the first prompt learning block, its input is the RGB tokens $\bm{T}_{\mathrm{RGB}}$ and the D-TIR fused representation $\bm{T}_{\mathrm{D-TIR}}$. For more detailed information of the prompt learning block, readers are referred to the reference (Zhu et al., 2023a).
@@ 5.1 Experimental Settings
ATOM Danelljan et al. (2019) RGB 0.649 0.695 2019
DiMP Bhat et al. (2019) RGB 0.662 0.710 2019
PrDiMP Danelljan et al. (2020) RGB 0.676 0.698 2020
ToMP Mayer et al. (2022) RGB 0.710 0.747 2022
OSTrack Ye et al. (2022) RGB 0.694 0.736 2022
SeqTrack Chen et al. (2023) RGB 0.719 0.766 2023
MixFormer Cui et al. (2024) RGB 0.732 0.781 2024
DeT_ATOM Yan et al. (2021b) RGB+D 0.639 0.665 2021
DeT_DiMP Yan et al. (2021b) RGB+D 0.667 0.700 2021
SPT Zhu et al. (2023b) RGB+D 0.706 0.757 2023
ViPT_RGBD Zhu et al. (2023a) RGB+D 0.720 0.759 2023
SDSTrack_RGBD Hou et al. (2024) RGB+D 0.718 0.763 2024
UnTrack_RGBD Wu et al. (2024) RGB+D 0.733 0.776 2024
TBSI Hui et al. (2023) RGB+T 0.690 0.749 2023
ViPT_RGBT Zhu et al. (2023a) RGB+T 0.693 0.752 2023
SDSTrack_RGBT Hou et al. (2024) RGB+T 0.666 0.708 2024
BAT Cao et al. (2024) RGB+T 0.713 0.782 2024
UnTrack_RGBT Wu et al. (2024) RGB+T 0.732 0.790 2024
DCEvo+OSTrack Liu et al. (2025) RGB+T 0.706 0.741 2025